# CAMeleon: Interactively Exploring Craft Workflows in CAD

### v1

@@ Abstract.
Designers of physical objects make assumptions on the material and fabrication workflow early in the design process. Recovering from bad assumptions is hard, because the design and resulting CAD model are locked-in to those assumptions. We present CAMeleon, a software tool to interactively explore fabrication workflows at any stage of the CAD process.
CAMeleon’s modular architecture allows users to execute their design with different workflows, and preview results. Users can freely explore alternative workflows. CAMeleon’s architecture, can be extended with new workflows, increasing the richness of workflows available.
We implemented five fabrication workflows in CAMeleon. We demonstrate CAMeleon’s extensibility through collaboration with six craftsmen whose workflows we replicated. We also implemented workflows of three papers and reflect on the process of these nine extensions. A usability study (n=12) showed that CAMeleon allowed participants to explore and select workflows for their design which they did not know before.
@@ 1. INTRODUCTION
Advances in computer-aided design (CAD) software, and computational fabrication techniques (e.g., additive manufacturing, subtractive fabrication, and wire forming) have enabled an increasingly broad set of people to engage in the design of physical objects. This trend towards personal fabrication (Baudisch et al., 2017) enabled marginalized communities to create custom assistive technology (Hurst and Tobias, 2011; Parry-Hill et al., 2017; Higgins et al., 2022), allowed high-schools to make fabrication part of their curriculum (Baudisch et al., 2019; Roumen et al., 2022; Hartikainen et al., 2024), and empowered several exciting hardware start-ups (e.g., Prusa3D, bambu, and XTOOL). As this trend continues, some envision fabrication machines to become as ubiquitous as kitchen appliances (Baudisch et al., 2023).
Designing physical objects presents challenges, particularly for people who are not well-versed in fabrication processes. Before even starting to design objects, designers are expected to make assumptions about the final fabrication process, as it has implications for how to design the structure supporting that process. While designers can bypass this challenge by obtaining shape models through 3D scanning, downloading, or through generative AI. The resulting models typically lack critical information about the fabrication processes needed to bring the designs to life, and changing the fabrication process to better fit the needs of the designer, is all but impossible.
@@
To address these challenges, HCI researchers have developed tools to make users really good at designing with a specific process/material in mind. Tools like Kyub (Baudisch et al., 2019) and FlatFitFab (McCrae et al., 2014a) are specialized for laser cutting, Bridging the gap (Dumas et al., 2014) and Stress relief (Stava et al., 2012) for 3D printing, and Plushie (Mori and Igarashi, 2007) for textile design. These tools have made computational fabrication accessible by reducing the required upfront domain knowledge of fabrication processes and materials. The price of using such software is that designers lock themselves into one specific fabrication process from the get-go (e.g., designing for laser cutting or for 3D printing). Ironically, as designers make progress on their design, they learn more about limitations/constraints imposed by the fabrication processes they chose, while at the same time it gets ever harder to explore alternatives.
@@
On the other end of the spectrum, there is the industrial workflow where engineers first design in CAD and then use so called ”Computer Aided Manufacturing” (CAM) to convert their design to machine-specific instructions. While CAM generates instructions for different machines, it still requires a design that is optimized for that specific fabrication process. These CAM workflows are heavily machine-oriented (e.g., they convert STLs to g-code for printing), and fail to acknowledge that most fabrication processes (especially craft workflows) involve manual processes.
@@
In this paper, we propose CAMeleon, a software tool to interactively explore craft workflows in Computer Aided Design.CAMeleon, allows users to defer the decision of what fabrication workflow and material to use, and lets users interactively explore alternatives until they start fabricating. It achieves this by encapsulating craft and fabrication workflows as modular components of the interface, each taking volumetric shape as input, converting that to workflow-specific geometry, and exporting instructions for both machines and humans. We call these instructions human-centered CAM workflows. As the target volumetric shape remains intact, users can freely explore alternative workflows before making final decisions. The modular architecture of CAMeleon makes it easily extensible, further increasing the richness of workflows available to end-users.
We identify three core use cases for CAMeleon: (1) during design: supporting designers to explore and preview different workflows for their product, (2) after design or using generative AI: serving as a tool after design, in a generative AI process to create shape models, or when users browse shape models, and then explore different manifestations thereof through CAMeleon. (3) after fabrication: re-fabricating existing objects with different fabrication workflows through 3D scanning.
We surveyed common craft workflows shared online, we classified these into categories, and from each of these categories we implemented one workflow in CAMeleon. To study the extensibility of CAMeleon, in collaboration with six craft experts we replicated their workflows in CAMeleon, we furthermore implemented the workflows of three research papers. We reflect on the process of extending CAMeleon through implementation and through follow-up interviews with the expert craft experts. In a usability study with 12 participants, we found that participants were more likely to explore different fabrication processes, allowing them to learn about fabrication and craft.
@@
We conclude that deferring the decision of fabrication processes, and making this decision interactive, allows designers to develop a more explorative attitude on workflows for computational fabrication. Especially designers with expertise that deviates from fabrication processes benefit from the ability to revise their decisions later on. As shape-based CAD models become more accessible, and tools like CAMeleon provide methods to capture human-centered CAM workflows, similar to recipes for cooking appliances, we reduce the barrier to entry for many more users to start taking advantage of their ability to create a physical reality that fits their needs.
2. WALK-THROUGH
@@
Figure 2 illustrates the process of Peter creating a small stool for his child using CAMeleon. (a) Peter begins by downloading a 3D volumetric model (here an STL file) from Thingiverse, an online repository of 3D models. He uploads this model into CAMeleon. (b) Unsure of which workflow best meets his needs, Peter starts by exploring the available options. He browses through some options, and becomes familiar with the various workflows and the machines required for fabrication, displayed as icons beneath each workflow. Peter has access to a laser cutter, a 3D printer, and a wire bender, which allows him to consider different approaches for creating the stool. Given the need for a load-bearing structure with a smooth finish for his child’s safety, Peter explores and compares workflows that meet his requirements, he drags the ’Mold making,’ ’Wire mesh sculpture,’ ’Thermal Wire Cutting for Foam Fabrication’, and ’Epoxy coating’ workflows to the view-port. After comparing his options, Peter decides that the ’Epoxy coating’ workflow is the most suitable. (c) He changed the layer height to 3 mm to fit the material he has in store. Once he is satisfied with his design, Peter exports the machine-instruction files (here SVG for laser cutting).
@@
After exporting, CAMeleon assists Peter step by step in executing the manual steps to build this stool. Figure 3 illustrates the interactive human-centered CAM workflow. (a) Peter already did the first step: laser cutting. (b) Peter then assembles and glues the cut pieces, fitting each layer together to form the structure. (c) Step 3: Peter completes the assembly by sanding the laser-cut structure to achieve a smooth finish. (d) Finally, Peter applies a layer of epoxy resin, carefully coating the entire structure for a sealed and polished finish.
@@ 3. CONTRIBUTIONS, BENEFITS, AND LIMITATIONS
We make three core contributions:
(1) The development of an interactive CAD tool, CAMeleon, that allows users to explore and compare fabrication workflows. We discuss the implementation and design decisions that went into the development of CAMeleon and release the code as an open source project for others to extend.
(2) The concept of human centered CAM workflows that include manual operations to capture realistic fabrication processes/workflows beyond mere machine instructions.
(3) The sample implementation of 14 fabrication workflows, including those derived from collaborations with craftsmen and recent research papers, to provide interactive fabrication guidance. We reflect on this extension process and what it takes for others to extend this further. This furthermore serves as a bridge for experts to distribute their workflows and processes across more users.
Limitations include that not every process can simply be converted into another, occasionally there are size or other technical requirements specific to a process that prevent conversion. Processes in CAMeleon include warnings to users for such requirements (e.g., ”features in this model are too small for PVC pipe bending, some details will get lost”) and its interactive preview gives them a sense of the expected outcome. Another limitation is that in the current state, it takes quite a bit of knowledge of CAMeleon’s code base and Three.JS to implement new workflows, in the future we plan to design a specific extension interface for CAMeleon to make this more accessible to craftsmen and people who invent new fabrication processes.
@@ 4.2. Design for Volumetric Fabrication
Fabricating volumetric shape often relies on various support structures, such as molds and jigs. Each serve different purposes such as facilitating shaping (e.g., plaster molds in ceramics, templates in woodworking), reinforcing structures (e.g., rebar molds in construction, wooden frameworks in shipbuilding).
@@ 4.3. Documenting Workflows
In disciplines such as industrial manufacturing (Derhamy et al., 2018), and business process management (Goo, 2010), there has been a long standing focus on documenting workflows and processes. Such documentation has been instrumental to increase efficiency (Rai et al., 2009), detect and minimize errors (Schiff and Bates, 2010), and support continuous process improvement (Tibeme et al., 2018). It furthermore facilitates the transfer of knowledge and ideas across teams or enterprises (Zhao et al., 2018). In craft communities such workflow knowledge often is transferred from one to another and fails to benefit from similar distribution.
Some research has begun to collect and integrate multiple fabrication workflows into a single platform, making significant strides in advancing digital fabrication. For example, Vespidae offers flexibility in controlling fabrication machines, allowing them to manipulate key parameters. It features modular components that users can freely combine to form new code definitions, facilitating a deeper exploration of the design space (Fossdal et al., 2023). Similarly, Tandem addresses the complexity of experimental digital fabrication workflows, successfully capturing and recording workflows across multiple software tools, including machine tuning and manual interventions (Tran O’Leary et al., 2024). Research such as Process Products captures iterative processes by embedding this information directly into objects, which helps designers communicate design evolution and aids manufacturers in understanding tool usage (Tseng and Tsai, 2015). Similarly, Product versus Process explores how design documentation is created and used by authors and readers on platforms like Instructables111https://www.instructables.com/ (Tseng and Resnick, 2014).
@@ 5. Surveying craft workflows
The goal of CAMeleon is to explore craft workflows. To understand this space, we explored processes shared in online communities. We reviewed 150 winning projects from the ’Workshop’ category on Instructables444https://www.instructables.com (sorted by ’I Made Its’). We excluded workflows centered on microprocessor-based products, purely 3D printed items without any additional workflow, and those focused solely on polishing, to maintain an emphasis on fabrication workflows. This left us with 56 projects. The appendix contains a list of all workflows used for this process with links to their instruction pages. We proceeded to cluster these into categories to better understand the design space. In Figure 4, the different processes are shown and divided into the following categories: Wire Forming, Interlocking Structure, Stacked Slice Construction, Guide Structure, Mold Casting. Two workflows did not fit this categorization, one of them was creating shapes with epoxy resin and the other building boats from wooden spars and frames.
To start off with a representative subset of fabrication processes based on this classification, we picked one representative sample of each category as what we call foundational processes. Figure 5 shows an overview of these: (a) Thermal Wire Cutting for Foam Fabrication, (b) Mold Casting, (c) Laser-Cut Interlocking Structures, (d) Wire Mesh Fabrication, and (e) Stacked Slice Construction. The variety of these processes not only gives users five interesting workflows to start out with, but they also serve as good templates as building blocks for others to produce workflows. By building human-centred CAM workflows for these processes, we allow users to convert their design to be fabricated with these workflows and generate both machine instructions and manual guidance.
@@ 6. CAMeleon
In this section we discuss design considerations and technical details that form the basis of CAMeleon. We discuss the interface design, the implementation, and demonstrate how we designed the system to support three different use cases.

### v2

# CAMeleon: Interactively Exploring Fabrication Workflows in CAD

@@ Abstract.
Designers of physical objects make assumptions on the material and fabrication workflow early in the design process. Recovering from bad assumptions is hard, because the design and resulting CAD model are locked-in to those assumptions. We present CAMeleon, a software tool to interactively explore and compare fabrication workflows at the end of the design process.
CAMeleon’s modular architecture allows users to execute their design with different workflows, and preview results. Users can explore alternative workflows. CAMeleon’s architecture is extensible with new workflows, increasing the scope of workflows available.
Based on a survey of 150 fabrication workflows shared online, we implemented five fabrication representative workflows in CAMeleon. To validate CAMeleon, we collaborated with six craftspeople, replicating their workflows to gain insights into their practices. Additionally, we demonstrated CAMeleon’s extensibility by implementing workflows from three research papers and reflecting on the extension process. Our design workshop (16 participants) highlights how working with CAMeleon unlocks a creative attitude of exploring workflows and opened up participants to workflows they had not previously imagined. We conclude that tools like CAMeleon have the potential to drastically expand designer’s ability to experiment with new materials and fabrication workflows.
@@ 1. INTRODUCTION
Advances in computer-aided design (CAD) software and computational fabrication (e.g., additive manufacturing, subtractive fabrication, and wire forming) have enabled an increasingly broad set of people to engage in the design of physical objects. This trend towards personal fabrication (Baudisch et al., 2017) enabled marginalized communities to create custom assistive technology (Hurst and Tobias, 2011; Parry-Hill et al., 2017; Higgins et al., 2022), allowed high-schools to make fabrication part of their curriculum (Baudisch et al., 2019; Roumen et al., 2022; Hartikainen et al., 2024), and empowered several exciting hardware start-ups (e.g., Prusa3D, BambuLab, and XTOOL). As this trend continues, some envision fabrication machines to become as ubiquitous as kitchen appliances (Baudisch et al., 2023).
However, designing physical objects presents challenges, particularly for people who are not well-versed in fabrication workflows. Before even starting to design objects, designers make assumptions about the final fabrication workflow (e.g., Laser cutting, 3D Printing, Vacuum forming …)—choices that determine how they will design the structure, select the material, operate the fabrication machine, and perform the appropriate post-processing. This is challenging because they may not know what workflows are available to them in the first place, or how choosing one workflow over another has implications for the final design.
@@
To address these challenges, HCI researchers have developed tools to assist users really good at designing with a specific process/material in mind. Tools like Kyub (Baudisch et al., 2019) and FlatFitFab (McCrae et al., 2014a) are specialized for laser cutting, Bridging the gap (Dumas et al., 2014) and Stress relief (Stava et al., 2012) for 3D printing, and Plushie (Mori and Igarashi, 2007) for textile design. These tools have made computational fabrication accessible by reducing the required upfront domain knowledge of fabrication processes and materials. The price of using such software is that designers lock themselves into one specific fabrication process from the get-go (e.g., designing for laser cutting or for 3D printing). Ironically, as designers make progress on their design, they learn more about limitations/constraints imposed by the fabrication processes they chose, while at the same time, it gets ever harder to explore alternatives. For example, for the bunny of Figure 1, it may be hard to achieve some details in the ears using wire bending which designers only find out when they get to that detail, or it may be better to laser cut an internal structure if the designer decides to fabricate at larger scale later on in the process. Recovering from these early assumptions is challenging because many of the parameters of the CAD model (wires, plates and joints, or volumes) are now specific to the initially chosen workflow.
@@
On the other end of the spectrum, there is the industrial workflow where engineers first design in CAD and then use so-called ”Computer Aided Manufacturing” (CAM) to convert their designs to machine-specific instructions. While CAM generates instructions for different machines, it still requires a design that is optimized for that specific fabrication process. These CAM instructions are heavily machine-oriented (e.g., they convert STLs to g-code for printing) and fail to acknowledge that most fabrication workflows involve manual processes (such as assembly, polishing, curing, etc.), which in practice highly influence the decision of what workflows to use.
@@
In this paper, we propose CAMeleon, a software tool to interactively explore craft workflows. CAMeleon, allows users to defer their final fabrication decision and explore and compare fabrication workflows until they start fabricating. It achieves this by encapsulating fabrication workflows as modular components of the interface, each taking volumetric shape as input, converting that to workflow-specific geometry, and exporting instructions for both machines and humans. We call these instructions human-centered CAM workflows. As the target volumetric shape remains unchanged, users can freely explore alternative workflows before making final decisions. The modular architecture of CAMeleon makes it easily extensible, further increasing the richness of workflows available to end-users.
We surveyed 150 fabrication workflows shared online, we classified these into categories, and from each of these categories we implemented one representative core workflow in CAMeleon. To validate CAMeleon, we collaborated with six craft experts, replicating their workflows to gain insights into their thinking. Additionally, to study CAMeleon’s extensibility, we implemented workflows from three research papers. In a design workshop with 16 participants, we found that participants use CAMeleon to develop an attitude to explore fabrication workflows, resulting in them making good decisions about their final workflow even if they had never used some of them before.
@@
We conclude that deferring the decision of fabrication workflows, and making this decision interactive, allows designers to develop an exploratory attitude on workflows for computational fabrication. Especially designers with expertise that deviates from specific fabrication workflows, benefit from the ability to revise their decisions later on. As shape-based CAD models become more accessible, and tools like CAMeleon provide methods to capture human-centered CAM workflows, similar to recipes for cooking appliances, we hope to take a small step towards reducing the barrier to entry for many more users to start taking advantage of their ability to create a physical reality that suits their needs.
2. WALK-THROUGH: exploring workflows to fabricate a stool
Figure 2 illustrates the process of Peter creating a small stool for his child using CAMeleon.
Peter is a designer who has some experience with 3D printing but not much with other fabrication workflows. Initially, he attempted to design a stool model for 3D printing following online tutorials—adjusting parameters such as thickening walls, reducing overhangs, and modifying infill settings to meet 3D printing requirements. However, after completing the print, Peter realized that his available 3D printing material lacks the inherent strength needed for the stool, making it unsuitable. Additionally, the surface finish did not meet his expectations.
Unsure how to proceed, Peter searches for alternative workflows. Platforms like Instructables provide useful tutorials and step-by-step instructions for specific projects, but Peter finds it difficult to adapt his design to those workflows. Furthermore, while many workflows seem promising, he struggles to visualize how the final fabricated results would look for his model. To compare workflows, he has to navigate between multiple web pages, gathering fragmented information on fabrication methods and repeatedly adjusting his design. This process is time-consuming and impractical.
Frustrated with this fragmented process, Peter turns to CAMeleon, which runs straight in his browser. CAMeleon allows Peter to explore and compare various workflows rendered as visualization of the expected outcomes.
@@ 2. WALK-THROUGH: exploring workflows to fabricate a stool
As shown in Figure 2a, Peter begins by downloading a 3D volumetric model (here an STL file) from Thingiverse, a repository of 3D models. He uploads this model into CAMeleon. (b) Unsure of which workflow best meets his needs, Peter starts by exploring the available options. He browses through some options and becomes familiar with the various workflows and the machines required for fabrication, displayed as icons beneath each workflow. Peter has access to a laser cutter, a 3D printer, and a wire bender, which allows him to consider different approaches for creating the stool. Given the need for a strong material with a smooth finish for his child’s safety, Peter explores and compares workflows that meet his requirements, he drags the ’Mold making,’ ’Wire mesh sculpture,’ ’Thermal Wire Cutting for Foam Fabrication’, and ’Epoxy coating’ workflows to the view-port. After comparing the resulting visualizations, as well as the manual and machine-based manufacturing processes involved, Peter decides that the ’Epoxy coating’ workflow is the most suitable. (c) He changed the layer height to 3 mm to fit the material he has in store. Once he is satisfied with his design, Peter exports the machine-instruction files (here SVG for laser cutting).
After exporting, CAMeleon assists Peter step by step in executing the manual steps to build this stool. Although the guidance serves as a general description, CAMeleon provides detailed instructions and parameters to ensure that users understand each step more deeply and can replicate the process effectively.
@@
Figure 3 illustrates the interactive human-centered CAM workflow. (a) Peter adjusts the provided standard guidance to fit his specific parameters and performs the laser cutting. (b) Peter then assembles and glues the pieces to form the structure. (c) Peter completes the assembly by sanding the laser-cut structure to achieve a smooth finish. While this step might be challenging for novice makers, CAMeleon provides tutorial links to guide users through the process. (d) Finally, Peter applies a layer of epoxy resin, carefully coating the entire structure for a sealed and polished finish.
@@ 3. CONTRIBUTIONS, BENEFITS, AND LIMITATIONS
This paper makes three core contributions:
(1) The development of an interactive design tool, CAMeleon, that allows users to explore and compare fabrication workflows. We discuss the implementation and design decisions that went into the development of CAMeleon and open source the code for others to extend.
(2) The concept of human-centered CAM workflows to include manual operations to capture realistic fabrication processes/workflows beyond mere machine instructions.
(3) The sample implementation of 14 fabrication workflows, including those derived from collaborations with craftsmen and recent research papers, to provide interactive fabrication guidance.
We evaluate these contributions with expert evaluation to include new workflows, and a design workshop to study usability and the role of CAMeleon in a collaborative design process.
Limitations include that not every model is suitable for all workflows; occasionally, there are size or other technical requirements specific to a process that prevent conversion. Processes in CAMeleon include warnings to users about such requirements (e.g., when laser cutting: ”features in this model are smaller than the chosen material thickness, you may experience a loss in fidelity”), and its interactive preview provides users with a sense of the expected outcome to sanity-check results before converging on a given workflow. In its current form, CAMeleon assumes one fabrication workflow for the entire model, while expert users can break down models into different sub-assemblies to then use different workflows (as evidenced by a participant in our workshop), there is currently no inherent support for this. Another limitation is that, in its current state, implementing new workflows requires knowledge of CAMeleon’s code base and Three.JS. In the future, we plan to develop an interface to design extensions for CAMeleon to make it more accessible to add new fabrication processes.
@@ 4.2. Design for Volumetric Fabrication
Fabricating volumetric shapes often relies on various support structures, such as molds and jigs. Each serves different purposes, such as facilitating shaping (e.g., plaster molds in ceramics, templates in woodworking) and reinforcing structures (e.g., rebar molds in construction, wooden frameworks in shipbuilding) or ensuring dimensional accuracy.
@@ 4.3. Documenting Workflows
In disciplines such as industrial manufacturing (Derhamy et al., 2018), and business process management (Goo, 2010), there has been a long standing focus on documenting workflows and processes. Such documentation has been instrumental to increase efficiency (Rai et al., 2009), detect and minimize errors (Schiff and Bates, 2010), and support continuous process improvement (Tibeme et al., 2018). It furthermore facilitates the transfer of knowledge and ideas across teams or enterprises (Zhao et al., 2018). In craft communities such workflow knowledge often is not transferred from one to another and fails to benefit from similar distribution due to unpredictable material behaviors (Ingold, 2010) and domain-specific tacit knowledge (Sennett, 2008).
To tackle this challenge, researchers have started integrating multiple fabrication workflows into unified platforms, significantly advancing the field of digital fabrication. For example, Vespidae offers flexibility in controlling fabrication machines, allowing them to manipulate key parameters. It features modular components that users can freely combine to form new code definitions, facilitating a deeper exploration of the design space (Fossdal et al., 2023). Similarly, Tandem addresses the complexity of experimental digital fabrication workflows, successfully capturing and recording workflows across multiple software tools, including machine tuning and manual interventions (Tran O’Leary et al., 2024). Research such as Process Products captures iterative processes by embedding this information directly into objects, which helps designers communicate design evolution and aids manufacturers in understanding tool usage (Tseng and Tsai, 2015). Similarly, Product versus Process explores how design documentation is created and used by authors and readers on platforms like Instructables111https://www.instructables.com/ (Tseng and Resnick, 2014).
@@ 5. Surveying craft workflows
The goal of CAMeleon is to let users explore fabrication workflows. To better understand the space of existing workflows, we explored workflows shared in online communities. We reviewed 150 winning projects from the ’Workshop’ category on Instructables444https://www.instructables.com (sorted by ’I Made Its’). We excluded workflows centered on microprocessor-based products, purely 3D printed items without any additional workflow, and those focused solely on polishing, to maintain an emphasis on fabrication workflows. This left us with 55 projects. The appendix contains a list of all workflows used for this process with links to their instruction pages. We proceeded to cluster these into categories to better understand the design space. In Figure 4, the different processes are shown and divided into the following categories: Mold Casting, Stacked Slice Construction, Interlocking Structure, Guide Structure, and Wire Forming. Two workflows did not fit this categorization; one of them was creating shapes with epoxy resin, and the other was building boats from wooden spars and frames.
To start off with a representative subset of fabrication processes based on this classification, we implemented one representative sample of each category in CAMeleon, we call these representative workflows. Figure 5 shows an overview of these at the example of the Stanford Bunny:(a) Thermal Wire Cutting Foam, (b) Mold Casting, (c) Laser-Cut Interlocking Structures, (d) Wire Mesh Fabrication, and (e) Stacked Slice Construction. The variety of these processes not only gives users five interesting workflows to start out with, but they also serve as good templates as building blocks for others to produce workflows.
@@ 6. CAMeleon
CAMeleon is a platform to explore and compare fabrication workflows. CAMeleon aims to support the process of computational fabrication and human-centered CAM workflows. We have the following design goals:
Shared and Reusable. CAMeleon’s code base should be modular by exposing workflow functions as reusable modules to enable extensibility.
Accessible Interface Design. The goal is to give users all the information needed to compare and explore workflows in a nutshell. Therefore, we focus on ease of use and good overviews. The hope is that designers at various levels of expertise are able to come to a useful comparison/decision.
Inclusive Open Platform. To bring together designers with different backgrounds, CAMeleon should not depend on existing software, and therefore should be implemented as browser tool. This furthermore ensures broad compatibility across devices and operating systems.
Compatible with Machines. To maximize support across fabrication machines, CAMeleon should support exporting to popular extensions for 3D (STL), 2D (SVG) and 1D/other (CSV) machines.
In this section, we discuss design considerations and technical details that form the basis of CAMeleon. We discuss the interface design and the implementation.